STIMULATED EMISSION PUMPING (SEP) SPECTROSCOPY APPLIED TO THE METHOXY RADICAL

JINJUN LIU, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210 (present address: Laboratory of Physical Chemistry, ETH Zurich, Building HCI, Wolfgang-Pauli-Str. 10, 8093 Zurich, Switzerland); \underline{MING-WEI CHEN}, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210; JOHN T. YI, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210 (present address: Department of Chemistry, Winston-Salem State University, Wilveria B. Atkinson Science Bldg, Room 311, Winston-Salem NC 27110); TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

Methoxy radical, CH_{3}O is a widely studied and benchmark molecule, both for its spectroscopy and its role in chemical reactions. It possesses a ground ^{2}E state which is distorted by the Jahn-Teller interaction and split into ^{2}E_{1/2} and ^{2}E_{3/2} components by the spin-orbit interaction. The LIF spectra of the Ã^{2}A_{1}–X̃^{2}E transitions is well-known, but under jet expansion conditions the E_{1/2} component is not observed because it is ∼60cm^{−1} higher than E_{3/2} and not populated. A feasible way to study the features of the X̃^{2}E_{1/2} level is to use the Stimulated Emission Pumping (SEP) technique. We have combined our high-resolution laser-induced fluorescence (LIF) spectroscopic (Δν ∼200MHz) with a moderateresolution laser (Δν ∼0.2cm^{−1}). These lasers are controlled by a computer program, which permits both the pump and dump lasers to be fired at specified delays after the photolysis laser producing CH_{3}O. SEP spectra of CH_{3}O were recorded with a resolution of ≈300MHz linewidth and measured with a precision <100MHz and these data were included in a global data (LIF, SEP, microwave) fitting to determine the parameters of the X̃^{2}E_{1/2} state. The previous assignment of parity for the ^{2}E_{1/2} state is modified and the value of the spin-orbit splitting revised.